If the average energy released per fission in a reactor is 150MeV, determine the amount of radioactive sample consumed in every second such that tractor releases thermal energy at a rate of 359MW. 


1
Expert's answer
2021-12-30T15:21:04-0500

"W=\frac{NE}t,\implies"

"N=\frac{Wt}E=1.5\cdot 10^{19}."